A rigid tank contains 1.60 moles of an ideal gas. Determine the number of moles of gas that must be withdrawn from the tank to lower the pressure of the gas from 20.0 atm to 5.00atm. Assume the volume of the tank and the temperature of the gas remain constant during this operation.
Answer #1

P1 / N1 = P2 / N2

N2 = (5.00 / 20.0) * 1.60

= 0.4 moles

number of moles of gas that must be withdrawn = 1.6 - 0.4

= 1.2 moles